OpenStudy (anonymous):

if 100 = 2 and 500 = 1 then 300 = ???

OpenStudy (ahsome):

That question makes no sense \[100\neq2\]\[500\neq1\]

OpenStudy (gorv):

take their mean

OpenStudy (gorv):

Q said if ....

OpenStudy (ahsome):

Oh. Then I am not sure, cause at that point mathematics kinda falls into itself ;)

OpenStudy (gorv):

\[\frac{ 500+100 }{ 2 }=\frac{ 1+2 }{ 2 }\]

OpenStudy (gorv):

300=3/2